Compute for the centripetal acceleration and force on a proton orbiting a single proton (Bohr model of a hydrogen atom). The electron is orbiting around the proton at a distance of 5 × 10¹¹ m and a speed of

2 × 10⁵m/s.


1
Expert's answer
2021-04-29T10:46:16-0400

The centripetal acceleration:


"a_c=\frac{v^2}{r}=\frac{(2·10^5)^2}{5·10^{11}}=0.08\text{ m/s}^2."